A New High-Resolution Image of Atomic Hydrogen in the M81 Group of Galaxies

M. S. Yun^{†,∗}, P. T. P. Ho^{†}, & K. Y. Lo^{‡}

Submitted to Nature Received

† Harvard-Smithsonian Center for Astrophysics, 60 Garden Street, Cambridge, MA 02138

∗ California Institute of Technology, Robinson 105-24, Pasadena, CA 91125

‡ Astronomy Department, University of Illinois, Urbana, IL 61801

INTRODUCTION

One of the most important discoveries of the Infrared Astronomical Satellite (IRAS) is a large population of galaxies with quasar-like luminosity. Massive bursts of star formation appear to be fueled by the mergers of gas-rich systems or interaction-induced influx of gas^{1−4}. Because sophisticated numerical models of galaxy interactions and sensitive spectral line observations at high spatial resolutions are still relatively new, the detailed comparisons of the theory and observations become possible only recently. Here we present a spectacular new view of the nearby M81 group in the emission line of neutral atomic hydrogen (HI), an excellent tracer of the total gaseous content of a galaxy well beyond the optical image^{5}. An unprecedented insight on the tidal interactions in this nearby group is obtained by combining the high angular resolution achieved with the Very Large Array (VLA) and the wide field of view using the mosaic technique. The complex of gaseous filaments threading M81, M82, and NGC3077, demonstrates the violent disruption of this system via tidal interactions. Such high sensitivity and high resolution studies will also shed light on the full extent of gaseous disks and dark halos.

Located only 3.6 Mpc (or 11 million light years) away^{6}, the “grand design” spiral galaxy M81 (NGC 3031) and its two small companions M82 (NGC 3034) and NGC 3077 form one of the nearest and best studied compact groups of galaxies. This system is a favorite testing ground for the density-wave theory and disk dynamics^{7}, active Seyfert nuclei^{8,9}, as well as starburst activities^{10}. Although there is little sign of tidal interactions in optical light (Figure 1), previous HI studies have suggested tidal interactions^{11−18}. Our new mosaic map confirms and extends our earlier results for M82 (YHL hereafter)^{18} and for M81 and NGC 3077 by van der Hulst^{14}, namely that the “cloud of HI” surrounding all three galaxies is dominated by filamentary tidal structures.

OBSERVATIONS AND ANALYSIS

With the VLA in the D configuration, four fields centered around M82 were observed (3 hrs per field) on January 24-25, 1990, and 8 additional fields including M81 and NGC3077 were observed (2 hrs per field) on April 29 and May 1, 1991, using all 27 telescopes. Quasars 3C 48 and 3C 286 were used for flux and bandpass calibration, while the VLA calibrator 0831+557 was observed periodically for short term gain and phase calibration. A total bandpass of 3.125 MHz with 64 spectral channels yielded a spectral resolution of 48.8 kHz (10.3 km sec^{−1}). Because of the large north-south velocity gradient (∼ 600 km sec^{−1}) across the mapped region, velocity centers for different fields were shifted accordingly in order to map all the HI emission. Phase self-calibration^{19} improved the dynamic range of the images by a factor of 3. The continuum emission is subtracted in uv-space using the average of line-free channel data.

NEUTRAL HYDROGEN GAS DISTRIBUTION

The HI distribution in the M81 group, shown in Figure 1, is the first HI image of the entire 150 kpc region at 1 kpc spatial resolution. The total neutral atomic hydrogen detected, 5.2 × 10^{9} M_{⊙}, is comparable to the HI mass for our Galaxy. Most remarkable is the extensive array of filamentary tidal features threading all three galaxies on a scale which has not been so clearly suggested in optical light. Each galaxy has a concentration of gas surrounding them, with the “South Tidal Bridge” connecting M81 and NGC 3077, and the “North Tidal Bridge” connecting NGC 3077 and M82. This is the first time that the North Tidal Bridge is traced in its full extent. The physical properties of the main HI features are summarized in Table 1.

GAS KINEMATICS

The gas kinematics traced in the intensity-weighted mean velocity map, Figure 2, indicates that the bulk of the gas around M81 partakes in the disk rotation while the remainder of the gas may be dominated by tidally induced motions. Kent^{20} finds that the mass of the large bulge and the disk alone is sufficient to reproduce the observed rotation out to a radius of 20 kpc (the only system in his sample of 16 galaxies which may not require any dark matter to explain the observed rotation). Our new extended rotation curve, Figure 3, shows a flattening near R ∼ 25 kpc, suggesting that a dark halo may be needed even for M81. The analysis of the disk kinematics becomes less accurate with increasing radius because of the tidally induced non-circular motion outside the 30 kpc radius and the presence of two massive companions M82 and NGC 3077 (each ∼ 3 × 10^{10} M_{⊙}) at projected distances of 40 kpc and 46 kpc. The gas kinematics along the South Tidal Bridge are clearly an extension of the outer disk rotation of M81, but the kinematics of the HI complex surrounding NGC 3077 and North Tidal Bridge seem to indicate their origin in NGC 3077 (see below). The tidal disruption of the HI disk of M82 suggested by YHL is further corroborated by the morphological and kinematic evidence in the new data.

THE MAJOR STRUCTURES

\underline{M81}. In optical light, M81 is a “grand design” two-arm spiral. In the inner disk, the HI spiral arms trace the optical arms very closely, but they bifurcate near the ends of the optical arms, and become much less tightly wound. The West Arm is difficult to trace in HI beyond the branching point, but the inner one appears to connect with Concentration I, the point where all disk features become severely disrupted. Both HI arms branching out of the East Arm continue on as two long spirals over 180^{∘} in angular extent, and the inner one continues onto the South Tidal Bridge. This association with a tidal feature and the non-circular, non-planar motions suggest that the HI arms are material arms rather than density wave driven. This is consistent with the bifurcated HI arms being outside the outer Lindblad resonance located at R ∼ 15 kpc^{21,25−26}. The outer HI arms are thus most likely tidally-driven transient spiral features^{24}. Each tidal encounter generally drive one pair of such transient arms. The double HI arms in M81 may be due to separate encounters with M82 and NGC3077. On the other hand, a tidal interaction can readily erase any pre-existing features, and the bifurcated HI arms may require other explanations.

Compared with the earlier HI studies of M81 by Gottesman & Weliachew^{21} and Rots & Shane^{22} (GW and RS hereafter), the HI distribution in our study is traced much further out (>30 kpc in radius) because of the multiple pointings with better sensitivity. “Concentrations I & II”, the HI complexes previously identified by GW as extragalactic, are now located inside the HI disk of M81. An exponential disk with a scale length of 10 kpc models the radial surface density distribution well out to 40 kpc radius. A significant excess found at R > ∼ 40 kpc where M82 and NGC 3077 are located, suggests that the HI of these galaxies dominates beyond this radius. There appears to be additional HI features located outside the mapped area, i.e. the western most feature of the M81 disk is the tip of another HI streamer pointing southwest^{23}.

\underline{“High Velocity Trough” and Ho IX.} The most significant deviation from the differential disk rotation in M81 is a ∼ 20 kpc long strip along the eastern border of the East Arm. This is seen as the Ω-shaped twists in the iso-velocity contours in Figure 2, with redshifts along the minor axis of up to 200 km sec^{−1} with respect to the systemic velocity of M81 (V = −15 km sec^{−1}). We call this the “High Velocity Trough (HVT)”. The associated HI mass is 3.3 × 10^{8} M_{⊙}, mostly in Concentration I, and the associated (lower limit) energy and momentum are 7.4 × 10^{55} ergs and 5.0 × 10^{10} M_{⊙} km sec^{−1}. Differential stretching of “supershells” or “superbubbles” formed by multiple supernovae or giant HII regions^{27} may offer an explanation, but the energy and momentum they provide are 10-1000 times too small. Instead, as in M101^{28}, the snowplow effect of the Magellanic type dwarf Ho IX (UGC 5336, DDO 66) found ∼ 10 kpc (in projection) east of M81, could explain the pile-up of gas near Concentration I and highly redshifted HI gas, if it passed in the plane of M81. However, empirical evidence^{29} suggest that such satellite systems generally occur away from the primary plane of disk galaxies.

A more likely alternative may be that these are the remnants of the recent collision between M81 and M82. The velocity of M82 matches the velocity perturbation associated with HVT both in sign and magnitude! The disk orbital period at 15 kpc radius is 5 × 10^{8} years, and the extent of the HVT suggests that M82 passed through the western disk of M81 near R ∼ 15 kpc some 2 × 10^{8} years ago. In this scenario, Ho IX, Concentration I, and the 10^{6} M_{⊙} molecular complex found near the HI peak^{30} can be explained as the resultant stellar and gaseous debris.

\underline{Arp Ring and Concentration II.} The main body of the Arp Ring^{31}, a faint halo-like optical structure found just north of M81, in fact coincides with the outer HI arm of M81. Concentration II is the brightest HI complex in the tidal bridge between M81 and M82 (Fig. 2), and it coincides with the brightest optical segment of Arp Ring. It is also the site where the tidally stripped gas from M81, the tidal streamer from M82, and North Tidal Bridge intersect, each with distinct velocities and velocity widths. The high integrated HI column density measured in the region may be due to the fortuitous superposition of several HI features seen along the line-of-sight. There is no kinematical evidence for a selfgravitating system. It appears that Arp Ring may not be a true feature but an ensemble of several unrelated tidal features around M81.

\underline{NGC 3077 and Tidal Bridges.} The North and South Tidal Bridges could be interpreted as a single continuous 100 kpc long tidal feature, which connects smoothly to one of the tidally induced spiral arms of M81. The observed velocity gradient is certainly consistent with an extension of the disk rotation of M81. However, the combined HI mass of 9 × 10^{8} M_{⊙} is nearly 1/3 of the total HI mass of M81, and would require a severe tidal disruption by a comparable or more massive companion over some 10^{9} years. No such companion exists, and NGC 3077 is certainly not massive enough. It has an early-type stellar continuum with little detectable velocity gradient across the stellar system and was thought to be a gas-poor dwarf elliptical^{32−34}. A giant molecular complex (M_{H_{2}} ∼ 10^{7} M_{⊙}) hosting vigorous star forming activity is found recently in the nuclear region^{35,36}. We find a local HI peak near the stellar nucleus while the bulk of the 7×10^{8} M_{⊙} HI complex is displaced by ∼ 4 kpc from the stellar peak. The HI velocity gradient across the major axis of NGC 3077 is consistent with rotation. The iso-velocity contours along South Tidal Bridge are perpendicular to its length as expected for a tidal tail (e.g., NGC 7252^{37}), but those along North Tidal Bridge, particularly near NGC 3077, are along the length, as might be expected for tidally induced shearing motion, similar to that of the remnant disk around M82. This suggests that NGC 3077 could be the origin for the gas complex around NGC 3077 and the North Tidal Bridge.

\underline{M82.} Earlier low resolution studies by Cottrell^{11} and others had interpreted the northsouth overall velocity gradient perpendicular to the stellar disk as that of a tidally captured HI cloud in a polar orbit. Our recent high resolution HI study of M82 using the VLA^{18} found luminous 10-20 kpc long HI streamers emerging from the warped stellar disk, and both the two-arm morphology and the smooth connection of their velocity fields onto the disk rotation are cited as the evidence for the tidal disruption of the outer HI disk of M82 by a massive companion (M81) in a polar orbit. The new study strengthens these conclusions by tracing previously discovered HI tails to greater and fainter extents (>30 kpc). A previously unidentified HI clump east of M82 (“Eastern Clump”) is now shown to be the anchoring base for the outer HI filaments. The large HI mass (8 × 10^{8} M_{⊙}) and sizes of the tidal tails suggest that the original HI disk of M82 was quite substantial prior to the tidal disruption by M81. Both the kinematics and the morphology suggest that the bulk of the material (M_{HI} = 8 × 10^{8} M_{⊙}, M_{H_{2}} = 2 × 10^{9} M_{⊙})^{38} is intrinsic to M82 rather than being captured from M81.

\underline{HI Clouds.} This experiment should have detected all the HI complexes more massive than 10^{6} M_{⊙} if their velocities fell within the observed range. Numerous HI clumps of mass 10^{6−7} M_{⊙} and ∼1 kpc diameter are seen in the channel maps, amid more diffuse emission of the HI disks or the tidal bridges, suggesting that they are local HI clouds or cloud complexes. No HI is detected for A0951+68, a dwarf spheroidal found ∼ 30′ south of M81. A possible HI satellite system is found as a bright and compact HI clump near α(1950) = 09^{h} 55^{m} 24^{s} and δ(1950) = +68^{∘} 53′ 48′′. It stands out from the overall HI distribution and kinematics – redshifted by 100 km sec^{−1} with respect to the disk rotation and tidal motion. Its centroid velocity V_{LSR} = +17 km sec^{−1} is similar to that of Galactic emission, but its large line width (ΔV_{F W ZI} ∼ 40 km sec^{−1}) suggests its extragalactic nature. If at the same distance as M81, its HI mass is 1.5 × 10^{6} M_{⊙} with a diameter of 1.3 kpc.

INDUCED ACTIVITIES AND EXTENT OF GALAXIES

This study demonstrates the significance of tidal interactions on the evolution of galaxies in small groups, even when there is little sign of close interactions in optical images. That the larger companion M81 seems less affected by the tidal interactions, may be due to its larger mass. Throughout this system, the outer disk gas which is less tightly bound gravitationally, appears more susceptable to tidal interactions. Note that the HI extent of M81 group is nearly 150 kpc in diameter. Maloney^{39} inferred a significantly larger extent (360-590 kpc radius) for disk galaxies from the large number of low redshift Lyman-α absorption systems found in the local universe. The observations presented here suggest that the presence of even a minor companion may drastically increase the apparent gaseous extent of galaxies. An HI absorption experiment against a nearby background quasar found HI gas at a projected distance of 65 kpc southeast of M81, outside the lowest contour in Figure 1, at the velocities predicted by our tidal analysis (D. Womble, private communication).

The new extended HI rotation curve for M81 suggests that a dark halo may dominate the dynamics outside the 20 kpc radius with an inferred mass of 2.1 × 10^{11} M_{⊙} inside the 30 kpc radius. If M82 is dynamically bound to M81, an even larger dynamical mass of about 10^{12} M_{⊙} is implied for M81 inside a 60 kpc radius. (The systemic velocity of NGC 3077 is nearly identical to M81 and provides little useful information). Similarly large and massive halos are also inferred from an ensemble of satellites orbiting disk galaxies by Zaritsky et al.^{40}. Very extended halos (100’s kpc diameter, more than typically modelled by the self-consistent models) such as suggested by White & Zaritsky^{41}, may offer a plausible explanation for both the flat rotation curve and the cores of the interacting galaxies mimicking two-body interactions at 40 kpc scale.

REFERENCES

1. Sanders, D.B., Soifer, B.T., Elias, J.H., Madore, B.F., Matthews, K., Neugebauer, G., & Scoville, N.Z. Astrophys. J. 325, 74-91 (1988).

2. Scoville, N.Z., & Soifer, B.T. Massive Stars and Starburst, (ed. K. Leitherer) 233-252 (Cambridge Univ. Press, Cambridge, 1991).

3. Barnes, J.E. Astrophys. J. 331, 669-717 (1988).

4. Barnes, J.E., & Hernquist, L.E. Astrophys. J. 370, L65-L69 (1991).

5. Briggs, F. H., Wolfe, A. M., Krumm, N., & Salpeter, E. E. Astrophys. J. 238, 510-523 (1980).

6. Freedman, W.L., Hughes, S.M., Madore, B.F., Mould, J.R., Lee, M.G., Stetson, P.B., Kennicutt, R.C., Turner, A., Ferrarese, L., Ford, H., Graham, J.A., Hill, R., Hoessel, J.G., Huchra, J., & Illingworth, G.D. Astrophys. J. 427, 628-655 (1994).

7. Visser, H. C. D. Structure and Properties of Nearby Galaxies (eds. E. M. Berkhuijsen & R. Wielebinski) 105-112 (Reidel, Dordrecht, 1978).

8. Peimbert, N., & Torres-Peimbert, S. Astrophys. J. 245, 845-856 (1981).

9. Wisniewski, W. Z., & Kleinmann, D. E. Astron. J. 73, 866-867 (1968).

10. Rieke, G. H., Lebofsky, M. J., Thompson, R. I., Low, F. J., & Tokunaga, A. T. Astrophys. J. 238, 24-40 (1980).

11. Roberts, M. S. External Galaxies and Quasi Stellar Objects (ed. D. S. Evans) 12-36 (Reidel, Dordrecht, 1972).

12. Cottrell, G. A. Mon. Not. R. astr. Soc. 178, 577-589 (1977).

13. Gottesman, S. T., & Weliachew, L. Astrophys. J. 211, 47-61 (1977).

14. van der Hulst, J. M. Astr. Astrophys. 75, 97-111 (1979).

15. Rots, A. H. Astr. Astrophys. Suppl. 41, 189-209 (1980).

16. Appleton, P. N., Davies, R. D., & Stephenson, R. J. Mon. Not. R. astr. Soc. 195, 327-352 (1981).

17. Cottrell, G. A. Mon. Not. R. astr. Soc. 174, 455-466 (1976).

18. Yun, M. S., Ho, P. T. P., & Lo, K. Y. Astrophys. J. 411, L17-L20 (1993).

19. Cornwell, T. J. 1982, Proceedings of the NRAO-VLA workshop, Vol. 13, 1 (NRAO, Green Bank, 1982).

20. Kent, S. M. Astron. J. 93, 816-832 (1987).

21. Gottesman, S. T., & Weliachew, L. Astrophys. J. 195, 23-45 (1975).

22. Rots, A. H., & Shane, W. W. Astr. Astrophys. 45, 25-42 (1975).

23. Appleton, P. N., & van der Hulst, J. M. Mon. Not. R. astr. Soc. 234, 957-969 (1988).

24. Toomre, A. & Toomre, J. Astrophys. J. 178, 623-666 (1972).

25. Lin, C. C., Yuan, C., & Shu, F. H. Astrophys. J. 155, 721-746 (1969).

26. Rots, A. H. Astr. Astrophys. 45, 43-55 (1975).

27. Heiles, C. Astrophys. J. Suppl. Ser. 55, 585-595 (1984).

28. van der Hulst, T., & Sancisi, R. Astron. J. 95, 1354-1359 (1988).

29. Holmberg, E. Ark. Astron. 5, 305-343 (1969).

30. Brouillet, N., Henkel, C., & Baudry, A. Astr. Astrophys. 262, L5-L8 (1992).

31. Arp, H. Science 148, 363-364 (1965).

32. Demoulin, M. Astrophys. J. 157, 81-85 (1960).

33. Barbieri, C., Bertola, F., & di Tullio, G. Astr. Astrophys. 35, 463-466 (1974).

34. Price, J. S., & Gullixson, C. A. Astrophys. J. 337, 658-670 (1989).

35. Becker, R., Schilke, P., & Henkel, C. Astr. Astrophys. 211, L19-L22 (1989).

36. Thronson, H. A., Wilton, C., & Ksir, A. Mon. Not. R. astr. Soc. 252, 543-549 (1991).

37. Hibbard, J. E., Guhathakurta, P., van Gorkom, J. H., & Schweizer, F. Astron. J. 107, 67-89 (1994).

38. Young, J. S., & Scoville, N. Z. Astrophys. J. 287, 153-166 (1984).

39. Maloney, P. Astrophys. J. 398, L89-L93 (1992).

40. Zaritsky, D., Smith, R., Frenk, C., & White, S.D.M. Astrophys. J. 405, 464-478 (1993).

41. White, S.D.M., & Zaritsky, D. Astrophys. J. 394, 1-6 (1992).

42. Brandt, J. C. Astrophys. J. 131, 293-303 (1960).

ACKNOWLEDGEMENTS. NRAO and VLA are operated by Associated Universities, Inc., under cooperative agreement with the National Science Foundation.

FIGURE LEGENDS

Fig. 1.— (Right) Red Palomar Sky Survey image of M81 triplet. Other than the optical filaments seen along the minor axis of M82, there is little hint of unusual activities in these galaxies in stellar light. (Left) Integrated HI map of M81 group from the new D-array VLA mosaic observations (θ_{F W HM} = 62′′ × 56′′). The mosaic of the data from 12 individual fields is made by noise-weighted linear averaging of the de-convolved (CLEANed) channel maps after correcting for the primary beam response. The inner 200 λ region of the uvspace is poorly sampled by the VLA, but nearly all the flux from single dish studies^{16} is recovered. The data cube contains a total of 81 channels covering the velocity range of −338 km sec^{−1} to +486 km sec^{−1}. By overlapping adjacent fields at half-beam spacings, we achieved uniform rms noise of 0.8 mJy beam^{−1} over the sampled region. Three channels centered at V_{LSR} = +12 km sec^{−1} and two channels near V_{LSR} = −50 km sec^{−1} contain Galactic HI emission: for the analysis of HI emission associated with the M81 group, these channels were clipped at the 10 mJy beam^{−1} level to remove the Galactic contribution. The linear grey scale represents the integrated HI flux. The contours correspond to the HI column density of 1.8 × 10^{20} cm^{−2} times 1, 2, 3, 4, 6, 10, 15, 25, and 40. Total of 5.2 × 10^{9} M_{⊙} of HI is detected.

Fig. 2. — Intensity weighted mean velocity (first moment) map of HI in M81 group. Contours are at 20 km sec^{−1} interval, and negative velocities are marked in dotted lines. The unit of the velocity label is in km sec^{−1}. The iso-velocity contours looking like “spider arms” centered on M81 are the classical signatures of differential rotation. This plus the motion of the tidally disrupted HI disk in M82 accounts for the north-south velocity gradient of this system. The differential rotation for the M81 disk appears regular for the inner 20 kpc, but the twisting iso-velocity contours in the outer regions (e.g. South Tidal Bridge) suggest significant non-circular motions at larger radii.

Fig. 3. — Mass contribution for the rotation curve of M81. The new rotation curve is traced out to 40 kpc by fitting the velocity field for a 90′ × 50′ elliptical region centered on M81. (20′ × 4′ strip along the eastern border of East Arm, “High Velocity Trough”, deviates significantly from the disk rotation and is masked off for this modeling.) The best fit Brandt curve^{42} is shown with a dotted line. The rotation speed rises to a maximum value of 250 km sec^{−1} around R ∼ 6 kpc and declines roughly in Keplerian manner until R ∼ 25 kpc. The analysis by Kent ^{20} shown here suggests that the mass of the disk and bulge alone could explain the rotation curve out to 20 kpc radius, but the flattening in rotation seen at 25 kpc may require a dark halo for M81 as well. The inferred dynamical mass for M81 is 2.1 × 10^{11} M_{⊙} inside the 30 kpc radius.